# On the Age of Information in Single-Server Queues with Aged Updates

### v1

@@ Abstract
The Age of Information (AoI) is a performance metric that quantifies the freshness of data in systems where timely updates are critical. Most state-of-the-art methods typically assume that packets enter the monitored system with zero age, neglecting situations, such as those prevalent in multi-hop networks or distributed sensing, where packets experience prior delays. In this paper, the AoI is investigated when packets have a non-zero initial age. We derive an expression for the average AoI in this setting, showing that it equals the standard AoI plus a correction term involving the correlation between packet age and inter-departure times. When these variables are independent, the expression simplifies to an additive correction equal to the mean initial age. In cases where the dependency structure is unknown, we also establish lower and upper bounds for the correction term. We demonstrate the applicability of our approach across various queueing scenarios, such as forwarding, tandem, and retrial queues. Additionally, we explore the accuracy of the derived bounds on a tandem composed of several queues, a model that has not yet been analytically solved from an age perspective.
@@ 1 Introduction
The Age of Information (AoI) is a relatively new metric that measures the freshness of the knowledge we have about the status of a remote system. More specifically, the AoI of a source is defined as the time elapsed since the generation of the last successfully received packet containing information about that source. The emergence of AoI in the seminal paper [1], has triggered an interest in its analysis in the context of queueing theory, which has been notably motivated by the fact that policies that optimise classical performance metrics, such as throughput, delay or package-loss probability, do not necessarily minimise the AoI.
@@
Most of the research in AoI considers the case in which the age of the packets to be delivered is zero, which corresponds to packets entering the communication channel as soon as they are generated. Under this assumption, a recent but already large body of literature has analyzed for example the average AoI [2], optimal scheduling [3], game theoretic aspects [4], energy harvesting [5], and other specific metrics such as peak AoI [6], Age of Incorrect of Information [7] and Query of AoI [8]. The book [9] provides a comprehensive survey on the topic, and a repository of papers on AoI is maintained at [10].
@@
The main result provides an expression for the Average Age of Information (AAoI) when packets have a non-zero initial age ($\Delta^{A}$), relating it directly to the AAoI in the standard case where packets arrive with zero age ($\Delta^{0}$). Specifically, it is proven that $\Delta^{A}=\Delta^{0}+\lambda^{e}\mathbb{E}\left[Y_{n}A_{n-1}\right],$ where $\lambda^{e}$ is the effective arrival rate of packets, and $A_{n-1}$ and $Y_{n}$ refer to the sequences of packet arrival ages and interdeparture times, respectively. In the case the age process and inter-departure times are statistically independent (which, as shown in Section 4, may occur in various scenarios), the equation reduces to $\Delta^{A}=\Delta^{0}+\mathbb{E}\left[A_{n-1}\right].$ For more complex systems where the dependency structure is unknown or cannot be studied analytically, specific bounds can be established to determine the range within which the final value of $\Delta^{A}$ lies. The applicability of the main result is illustrated in various settings, including a forwarding queue, tandem queues and retrial queues. Additionally, the validity of interval estimation for $\Delta^{A}$ is tested in some of these scenarios, and its practical usefulness is highlighted in an application to a multi-queued tandem.
The rest of the paper is organised as follows. In Section 2.1, the system model is introduced and the age process with aged updates is characterised. In Section 3, the main result of this article is presented and demonstrated. In Section 4, some illustrative applications of the model are described. Finally, some final comments are given in Section 5 together with some suggestions for future research.
@@ 2.1 System model
Consider a communication system through which a destination node receives updates about a process of interest. At epochs $g_{n}$, information updates are generated at the source and transmitted to the destination as packets. However, unlike standard models, we assume that updates do not enter the system immediately upon generation. Instead, they arrive at the queueing system after incurring a certain delay.
Packets enter the queueing system at epochs $t_{n}$ according to an arrival process of rate $\lambda$ with i.i.d. inter-arrival times $X_{n}=t_{n}-t_{n-1}$. Each arriving packet has a corresponding initial age $A_{n}\geq 0$, defined as the time elapsed between the packet’s generation and its arrival at the system. The initial ages $\left\{A_{n}\right\}$ are modelled as a stationary and ergodic sequence of identically distributed, but not necessarily independent, variables. Importantly, we allow $A_{n}$ to statistically depend on the inter-arrival times $X_{n}$ and the inter-departure times $Y_{n}$. This is a key aspect that introduces additional complexity in the model. Still, it also reflects a wide variety of realistic scenarios, enhancing the usefulness and relevance of this framework. The stationarity assumption on $A_{n}$ ensures that the process remains bounded in expectation and avoids pathological cases of ever-increasing update ages.
Upon arrival, a packet is either discarded, queued for future processing, or served immediately, depending on the queue size, the packet management policy, and the system’s state. Packets that are not discarded enter the system and complete the service. These are referred to as effective arrivals, and the effective inter-arrival times are denoted by $X_{n}^{e}$. Once a packet enters the server, it stays in service for a $\mu$-rate exponential time $S_{n}\sim\textsl{exp}\left(\mu\right)$, independent of the inter-arrival time. The service completion times are denoted by $t^{\prime}_{n}$, and define the sequence of departures. The system time of packet $n$ is $T_{n}=t^{\prime}_{n}-t_{n}$, and the inter-departure time is $Y_{n}=t^{\prime}_{n}-t^{\prime}_{n-1}$.
@@ 2.2 Age of Information with aged updates
A graphical illustration of the AoI evolution in this model is shown in Figure 2. The process evolves piecewise linearly, with occasional jumps at delivery times $t^{\prime}_{n}$. While it still maintains the typical sawtooth shape characteristic of such functions, some updates may enter the system with a high initial age. At the time of delivery of these updates, the age at the destination may increase instead of resetting to a lower value. This behaviour is a distinctive feature of models with aged updates and is not seen in standard AoI models with zero-aged updates [11]. Specifically, at times $t^{\prime}_{n}$, the AoI resets from the peak value $A_{n-1}+X_{n}+T_{n}$ to $A_{n}+T_{n}$. Thus, if $A_{n}$ exceeds the threshold $X_{n}+A_{n-1}$, the delivery of this packet leads to an upper jump in the AoI function, as illustrated by packet 3 in Figure 2. We term these ”far updates” due to their similarity to distant status updates in asynchronous monitoring.
@@ 3 Main results
For the single-server queue with aged updates described in Section 2.1, the AAoI is given by
@@
where $\Delta^{0}$ is the AAoI in the corresponding system with zero-aged updates, $\lambda^{e}$ is the effective arrival rate, $Y_{n}$ is the $n$-th inter-departure time, and $A_{n-1}$ is the initial age of the $(n-1)$-th packet.
@@
The graphical method is adopted, a widely used technique for computing the AAoI via geometric area decomposition [6, 12, 11]. Let us examine Figure 2, which illustrates the evolution of the AoI in a system with aged updates, including a sample far update, and suppose that the process is observed during an interval $[0,T]$. Without loss of generality, it can be assumed that deliveries occur at both the initial time $t=0$ and the final time $t=T$, that is, $t^{\prime}_{0}=0$ and $t^{\prime}_{n}=T$ for some $n\leq N(T)$, where $N(T)$ is the number of delivered updates at the end of the observed period, as defined in (1). This is a harmless assumption since the only consequence of leaving it out is the addition of two finite areas that, nonetheless, will vanish in the limit. Also, for simplicity, assume that after the delivery at $t=0$, the system becomes empty.
@@
However, the integral in Equation (3) can be replaced by the sum of the disjoint polygonal areas $H_{n}$, for $n=0,1,\ldots,N(T)-1$. Each of these areas consists of the sum of a triangular area and a rectangular area,
@@
Hence, Equation (3) can be written as
@@
By the Global Balance Principle, the flow out of a set of states is equal to the flow into the same set. By applying this principle to the entire set of states, it follows that the effective arrival rate and the departure rate are equal. Additionally, the ratio between the number of delivered packets and the interval length approaches the departure rate. Thus,
$\lambda^{e}\stackrel{{\scriptstyle def}}{{=}}1/\mathbb{E}\left[X^{e}\right]=1/\mathbb{E}\left[Y\right]=\lim_{T\to\infty}\dfrac{N(T)}{T}.$
The limit in Equation (6) can then be expanded as
@@
The first two terms in the parentheses are recognised as those defining the standard AAoI with zero-age updates [11, 6], and the proof is $\qquad{}_{\square}$.
A well-known identity in probability says that the expected value of the product of two random variables can be expressed as the sum of their covariance and the product of their means. Moreover, the coefficient of variation of a random variable is defined as the ratio between the standard deviation and the mean, while the correlation coefficient between two random variables is the ratio between their covariance and the product of the standard deviations. Because of this, the age correction term in Theorem 1 can be rewritten as
@@
$=\mathbb{E}\left[A_{n}\right]+r_{Y,A}\cdot\kappa_{Y}\cdot\sigma_{A},$ (7)
where $r_{Y,A}$ is the correlation coefficient between $Y_{n}$ and $A_{n-1}$, and $\sigma_{*}$ and $\kappa_{*}$ stand, respectively, for the standard deviation and the coefficient of variation of the corresponding random variable.
@@
The above expression states that, given the marginal distributions of $Y_{n}$ and $A_{n-1}$, the dependence between them affects the correction term only through the correlation coefficient. From another perspective, (7) can also be interpreted as implicitly limiting the correlation between $Y_{n}$ and $A_{n-1}$. Indeed, since the correction term is positive, the correlation coefficient must satisfy
$\mathbb{E}\left[A_{n}\right]+r_{Y,A}\cdot\kappa_{Y}\cdot\sigma_{A}>0\ \Longrightarrow\ r_{Y,A}>-\kappa_{A}^{-1}\kappa_{Y}^{-1}.$
@@
The following corollary is also a direct consequence of equation (7).
@@ 4.1 Independency
The model analysis greatly simplifies when the sequence $\{A_{n}\}$ of initial ages is independent of the sequence of inter-departure times $\{Y_{n}\}$. In this case, the cross term of Theorem 1 factorizes as $\mathbb{E}\left[Y_{n}A_{n-1}\right]=\mathbb{E}\left[Y_{n}\right]\mathbb{E}\left[A_{n-1}\right]$, and the AAoI can be then expressed as
@@
This case is relevant in real-world applications where sources and monitors are not perfectly synchronised. For instance, [13] reports observations of fixed delays in practical communication systems due to clock shifts. Our framework naturally incorporates such delays, allowing for a direct and simple quantification of their impact on information freshness.
@@ 4.1.1 Heavy traffic
where $x^{+}=\max\{0,x\}$, defines the inter-departure times as the sum of the service and the idle times. As a result, under HT conditions, the inter-departure times asymptotically converge in distribution to the service time distribution, and therefore become independent of the arrival process and initial ages. Consequently, in Equation (7), $r_{Y,A}\to 0$, and the correction term collapses to $\mathbb{E}\left[A_{n}\right]$. Furthermore, since $\kappa_{Y}\to 1$, the range of the interval defined in Corollary 1 approaches $d\to 2\sigma_{A}$, resulting in a tighter interval for stable age processes.

### v2

@@ Abstract
[…]
@@ 1 Introduction
The Age of Information (AoI) is a relatively new metric that measures the freshness of the knowledge we have about the status of a remote system. More specifically, the AoI of a process is defined as the time elapsed since the generation of the last successfully received packet containing information about that process. The emergence of AoI in the seminal paper [kaul:2011] has triggered an interest in its analysis in the context of queueing theory, notably motivated by the fact that policies that optimise classical performance metrics, such as throughput, delay or package-loss probability, do not necessarily minimise the AoI.
@@
Most of the research in AoI considers the case in which the age of the packets to be delivered is zero, which corresponds to packets entering the communication channel as soon as they are generated. Under this assumption, a rich body of literature has developed, providing a variety of analytical tools and foundational results. This includes the foundational analysis of the average AoI in various queueing systems using methods such as the graphical approach [kaul_real:2012, costa:2016, yates:2019, soysal:2021, lin:2022] and Stochastic Hybrid Systems (SHS) [yates:2018, yates:2019, yates:2020, kam:2022], optimal scheduling policies aimed at minimising AoI [bedewy:2019, yates:2019], and investigations into AoI behaviour under energy harvesting constraints [arafa:2019]. The field has also expanded beyond just the average AoI to explore a family of age-related metrics, such as the peak AoI [costa:2016], the Age of Incorrect Information [maatouk:2020], and the Query AoI [chiariotti:2022]. Additionally, researchers have examined these concepts through game-theoretic perspectives [nguyen:2018, miguelez:2023]. More recent and general analytical efforts have focused on deriving the full probability distribution of the AoI [inoue:2019, yates:2020] and establishing performance bounds for a wider range of systems [farazi:2019, soysal:2021]. Comprehensive surveys [sun:2020, yates:2021] and texts [kosta:2017] document the rapid expansion of this field. A repository of papers related to AoI can be found at [AoIRepository].
@@
The analysis of AoI in multi-stage systems is gaining significant attention, with recent works making substantial progress in specific directions. These include the quantitative characterisation of AAoI in tandems of queues [kam:2022], developing scheduling policies for multi-hop wireless networks [talak:2017], characterising age accumulation under preemptive service disciplines [yates:2018], deriving bounds for systems with cross-traffic [vikhrova:2022], and exploring advanced concepts like optimisation under imperfect knowledge [zhao:2025], peak-age analysis in cache-enabled networks [ernest:2025], and semantics-aware status updates [delfani:2025].
However, a fundamental challenge remains: the conventional approach of modelling each multi-stage system from scratch often leads to highly specific and potentially intractable derivations. As a result, there is a scarcity of general analytical results that can be applied across different system configurations, as acknowledged in [farazi:2019]. The proposal presented in this paper can therefore be seen as a novel view to address this challenge. Instead of analysing the entire system complexity at once, we model the final communication link as a single-server queue where packets arrive with a non-zero initial age, effectively decoupling the analysis of the prior system’s impact from the dynamics of the final queue.
In our main contribution, we provide an expression for the Average Age of Information (AAoI) when packets have a non-zero initial age ($\Delta^{A}$), relating it directly to the AAoI in the standard case where packets arrive with zero age ($\Delta^{0}$). The resulting decomposition, given in Theorem 1, serves as a unified tool to tackle a wide class of problems, enabling both the recovery of known results and the derivation of new expressions and bounds for complex systems where traditional methods struggle.
It is important to note that the proposed framework is designed to be general and does not assume any specific packet management policy, such as discarding obsolete packets at the monitor. This choice is deliberate for two reasons. First, it allows the model to be applied to systems where such discarding is not feasible, such as those requiring in-order delivery for state consistency, complete data reconstruction, or non-volatile logging. Second, and more importantly, the effect of any packet management policy is inherently captured by our model through its impact on the effective arrival rate, $\lambda^{e}$, and the dependence structure between the initial ages and the departure process at the final queue. This versatility is a key strength of the proposed approach.
The main contributions of this work are:
  • General framework: We derive an expression for the AAoI in a single-server queue with aged updates (Theorem 1). The result shows that the AAoI equals the standard AAoI plus a correction term involving the correlation between the initial age and inter-departure times. When these variables are independent, the expression simplifies to an additive correction equal to the mean initial age.
  • Performance bounds: For cases where the dependency structure is unknown, we establish lower and upper bounds for the correction term (Corollary 1) that depend only on the marginal distributions of the involved processes.
  • Unification of results: We demonstrate that the framework can recover AAoI expressions for well-studied systems, such as forwarding queues and tandem queues, under a common analytical approach.
[…]
  • Practical utility: We explore the accuracy of the derived bounds in complex scenarios, such as a tandem of multiple queues, and validate their effectiveness through simulations.
The rest of the paper is organised as follows. In Section 2, the system model is introduced and the age process with aged updates is characterised. In Section 3, the main result of this article is presented and demonstrated. In Section 4, some illustrative applications of the model are described. Finally, some final comments are given in Section 5 together with some suggestions for future research.
@@ 2.1 System model
Consider a communication system through which a destination node receives updates about a process of interest. Information updates are generated at the source at epochs $g_{n}$ and transmitted to the destination as packets. Unlike standard models, these updates do not enter the queueing system immediately upon generation but arrive after incurring a certain delay.
Packets enter the queueing system at epochs $t_{n}$ according to an arrival process of rate $\lambda$ with i.i.d. inter-arrival times $X_{n}=t_{n}-t_{n-1}$. Each arriving packet has a corresponding initial age $A_{n}\geq 0$, defined as the time elapsed between the packet’s generation and its arrival at the system. The sequence of initial ages $\{A_{n}\}$ is assumed to be stationary and ergodic, with identically distributed (but not necessarily independent) random variables. Importantly, $A_{n}$ can be statistically dependent on the inter-arrival times $X_{n}$ and the future inter-departure times $Y_{n}$. This dependence is a key aspect of our framework, reflecting realistic scenarios and introducing additional analytical complexity. The stationarity assumption ensures the process remains well-behaved, avoiding cases of ever-increasing ages.
Upon arrival, a packet is either discarded, queued for future processing, or served immediately, depending on the queue size, the packet management policy, and the system’s state. Packets that are not discarded enter the system and complete the service. These are referred to as effective arrivals, and the effective inter-arrival times are denoted by $X_{n}^{e}$. Once a packet enters the server, it stays in service for a time $S_{n}\sim\textsl{exp}\left(\mu\right)$, independent of the arrival process. The service completion times are denoted by $t^{\prime}_{n}$, and define the sequence of departures. The system time of packet $n$ is $T_{n}=t^{\prime}_{n}-t_{n}$, and the inter-departure time is $Y_{n}=t^{\prime}_{n}-t^{\prime}_{n-1}$.
@@ 2.2 Age of Information with aged updates
A graphical illustration of the AoI evolution in this model is shown in Figure 2. The process evolves piecewise linearly, with occasional jumps at delivery times $t^{\prime}_{n}$. While it still maintains the typical sawtooth shape of age functions, some updates may enter the system with a high initial age, meaning they carry obsolete information. At the time of delivery of these updates, the age at the destination may increase instead of resetting to a lower value. This behaviour is a distinctive feature of models with aged updates and is not seen in standard AoI models with zero-aged updates [yates:2021]. Specifically, at times $t^{\prime}_{n}$, the AoI resets from the peak value $A_{n-1}+X_{n}+T_{n}$ to $A_{n}+T_{n}$. Thus, if $A_{n}$ exceeds the threshold $X_{n}+A_{n-1}$, the delivery of this packet leads to an upper jump in the AoI function, as illustrated by packet 3 in Figure 2. We term these ”far updates” due to their similarity to distant status updates in asynchronous monitoring.
The occurrence of far updates, while detrimental to age performance, is unavoidable in systems where packets cannot be discarded, such as those requiring in-order delivery (e.g., distributed consensus), data reconstruction (e.g., media streaming with error correction), or maintaining complete historical records (e.g., flight data recorders). Although average freshness may not be the primary concern in some of these applications, tracking the behaviour of the AoI can reveal latent flaws or performance bottlenecks in the system. The model is also relevant when AoI itself is the primary performance metric of the system, regardless of the information contained in the packet. In this context, discarding a high-age packet would artificially improve the metric and mask the true system performance. Furthermore, many systems naturally prevent the existence of far updates through their inherent dynamics, as demonstrated in Sections IV-C.1 and IV-C.3.
On the other hand, when packet discarding is implemented, its effect is inherently captured by our framework through the effective arrival rate $\lambda^{e}$ and the dependency structure between $A_{n}$ and $Y_{n}$. Thus, the proposed approach remains relevant regardless of the packet management policy.
@@ 3 Main results
For the single-server queue with aged updates described in Section 2, the AAoI is given by
@@
where $\Delta^{0}$ is the AAoI in the corresponding system with zero-aged updates, $\lambda^{e}\stackrel{{\scriptstyle def}}{{=}}1/\mathbb{E}\left[X^{e}\right]$ is the effective arrival rate (i.e., the rate of packets that are admitted into service), $Y_{n}$ is the $n$-th inter-departure time, and $A_{n-1}$ is the initial age of the (n1)(n-1)-th packet.
@@
The graphical method is adopted, a widely used technique for computing the AAoI via geometric area decomposition [costa:2016, yates:2019, yates:2021]. Let us examine Figure 2, which illustrates the evolution of the AoI in a system with aged updates, including a sample far update, and suppose that the process is observed during an interval $[0,T]$. Without loss of generality, it can be assumed that deliveries occur at both the initial time $t=0$ and the final time $t=T$, that is, $t^{\prime}_{0}=0$ and $t^{\prime}_{n}=T$ for some $n\leq N(T)$, where $N(T)$ is the number of delivered updates at the end of the observed period, as defined in (1). This is a harmless assumption since the only consequence of leaving it out is the addition of two finite areas that, nonetheless, will vanish in the limit. Also, for simplicity, assume that after the delivery at $t=0$, the system becomes empty.
@@
However, the integral in Equation 3 can be replaced by the sum of the disjoint polygonal areas $H_{n}$, for $n=0,1,\ldots,N(T)-1$. Each of these areas consists of the sum of a triangular area and a rectangular area,
@@
Hence, Equation 3 can be written as
@@
Under the stability condition $\mathbb{E}\left[S_{n}\right]<\mathbb{E}\left[X_{n}^{e}\right]$, the system reaches a steady state, and the sequence of effective inter-arrival times $\{X_{n}^{e}\}$ forms a stationary and ergodic process. By the principle of flow conservation [krakowski:1973], the long-term effective arrival rate equals the long-term departure rate. Additionally, the Elementary Renewal Theorem [ross:stochastic] asserts that the ratio between the number of delivered packets and the interval length approaches the departure rate. Therefore,
$\lambda^{e}=1/\mathbb{E}\left[X_{n}^{e}\right]=1/\mathbb{E}\left[Y_{n}\right]=\lim_{T\to\infty}\dfrac{N(T)}{T}.$
The limit in Equation 6 can then be expanded as
@@
The first two terms in the parentheses are recognised as those defining the standard AAoI with zero-age updates [yates:2021, costa:2016], and the proof is complete $\qquad{}_{\square}$.
Remark.
The framework naturally incorporates the effect of packet management policies, such as blocking or discarding updates at the monitor. The impact of any such policy is directly captured by two elements in Theorem 1:
  1. a)
    The effective arrival rate $\lambda^{e}$, which reflects only the packets admitted into the system. A discard policy would reduce $\lambda^{e}$ compared to the raw arrival rate $\lambda$.
  2. b)
    The dependence structure between $Y_{n}$ and $A_{n-1}$, as the policy determines which packets are served and thus defines the joint distribution of inter-departure times and initial ages.
Therefore, our general expression is also applicable in the case where any of these policies exists, providing a foundation for their analysis rather than being incompatible with them.
A well-known identity in probability says that the expected value of the product of two random variables can be expressed as the sum of their covariance and the product of their means: $\mathbb{E}\left[Y_{n}A_{n-1}\right]=\mathbb{C}\mathrm{ov}{}\left[Y_{n},A_{n-1}\right]+\mathbb{E}\left[Y_{n}\right]\mathbb{E}\left[A_{n}\right]$. Moreover, the coefficient of variation of a random variable is defined as the ratio between the standard deviation and the mean, $\kappa_{Y}=\sigma_{Y}/\mathbb{E}\left[Y_{n}\right]$, while the correlation coefficient between two random variables is the ratio between their covariance and the product of the standard deviations, $r_{Y,A}=\mathbb{C}\mathrm{ov}{}\left[Y_{n},A_{n-1}\right]/(\sigma_{Y}\sigma_{A})$, where $\sigma_{\cdot}$ stands for the standard deviation of the corresponding random variable. Because of this, the age correction term in Theorem 1 can be rewritten as
@@
$=\mathbb{E}\left[A_{n}\right]+r_{Y,A}\cdot\kappa_{Y}\cdot\sigma_{A}.$ (7)
@@
The above expression states that, given the marginal distributions of $Y_{n}$ and $A_{n-1}$, the dependence between them affects the correction term only through the correlation coefficient. From another perspective, Equation 7 can also be interpreted as implicitly limiting the correlation between $Y_{n}$ and $A_{n-1}$. To understand this, we first note that the correction term is always non-negative, because it is the product of a positive rate and the expected value of a non-negative random variable, $Y_{n}A_{n-1}$. Intuitively, the non-zero initial age $A_{n-1}$ represents a pre-existing delay that can only contribute to an increase in the overall age at the monitor, as it represents time that has already elapsed and cannot be “recovered”. The term $\mathbb{E}\left[Y_{n}A_{n-1}\right]$ quantifies the interaction between this prior delay and the system’s dynamics, always adding to the baseline age $\Delta^{0}$. Therefore, $\mathbb{E}\left[A_{n}\right]+r_{Y,A}\cdot\kappa_{Y}\cdot\sigma_{A}>0$, and the correlation coefficient must satisfy $r_{Y,A}>-\kappa_{A}^{-1}\kappa_{Y}^{-1}$.
@@
The following corollary is also a direct consequence of Equation 7.
@@ 4.1 Independency
We first examine the relevant special case in which the sequence $A_{n}$} of initial ages is independent of the sequence of inter-departure times $Y_{n}$}. Under this assumption, the model analysis greatly simplifies since the cross term of Theorem 1 factorises as $\mathbb{E}\left[Y_{n}A_{n-1}\right]=\mathbb{E}\left[Y_{n}\right]\mathbb{E}\left[A_{n-1}\right]$, and the AAoI can then be expressed as
@@
This case is relevant in real-world applications where sources and monitors are not perfectly synchronised. For instance, [beytur:2019] reports observations of fixed delays in practical communication systems due to clock shifts. Our framework naturally incorporates such delays, allowing for a direct and simple quantification of their impact on information freshness.
@@ 4.1.1 Heavy traffic
where $x^{+}=\max\{0,x\}$, defines the inter-departure times as the sum of the service and the idle times. As a result, under HT conditions, the inter-departure times asymptotically converge in distribution to the service time distribution, and therefore become independent of the arrival process and initial ages. Consequently, in Equation 7, $r_{Y,A}\to 0$, and the correction term collapses to $\mathbb{E}\left[A_{n}\right]$. Furthermore, since $\kappa_{Y}\to 1$, the range of the interval defined in Corollary 1 approaches $d\to 2\sigma_{A}$, resulting in a tighter interval for stable age processes.